How would I solve thisJoes Hardware is adding a new product line that will require an investment of $1,470,000. Managers estimate that this invest
Answer #1
Total cash flows 2555000 =310000+285000+(245000*8)
Less: Investment cost 1470000
Total net operating income 1085000
Average annual operating income 108500 =1085000/10
Average amount invested 735000 =1470000/2
Average annual operating income from Asset / Average amount invested = Accounting Rate of return
108500 / 735000 14.76%